AB-INITIO INVESTIGATION OF TERAHERTZ VIBRATIONAL MODES IN MOLECULAR CRYSTALS

\underline{PETER UHD JEPSEN}, Department of Photonics Engineering, Technical University of Denmark, DK-2800 Kongens Lyngby, Denmark; STEWART CLARK, Department of Physics, University of Durham, Durham DH1 3LE, United Kingdom.

We use the CASTEP ab-initio density-functional perturbation theory code^{a} to investigate the properties of terahertzfrequency vibrational modes in crystalline systems. We will discuss the strong mixing of molecular and phonon modes in hydrogen-bonded crystals^{b}, and also discuss ongoing work where we study nonlinear interaction between intense THz light and molecular crystals.

aS. J. Clark, et al., Z. Kristallographie 220, 567 (2005)

bP. Uhd Jepsen and S. J. Clark, Chem. Phys. Lett. 442, 275 (2007)